[figure: 9.7 cm]

LIEB-MATTIS MODEL FOR TWO INTERPENETRATING ANTIFERROMAGNETS

Dirk Schmalfuß, Sven E. Krüger, Johannes Richter Institut für Theoretische Physik Universität Magdeburg PF 4120, D-39016 Magdeburg, Germany

We investigate a model which can be considered as a reference model describing the unusual magnetic properties of Ba_{2}Cu_{3}O_{4}Cl_{2}. The main feature of this compound is the existence of two interpenetrating subsystems of antiferromagnetically coupled copper atoms, Cu(A) and Cu(B). The coupling between Cu(A) is strong while the coupling between Cu(B) is much weaker as indicated experimentally by two critical temperatures differing in one order of magnitude. An additional coupling between the subsystems gives rise to interesting frustration effects. Our approach is an exactly solvable spin-1/2 Heisenberg model with long-range Lieb-Mattis interactions. In the limit N → ∞ we get a complete thermodynamic phase diagram. We are able to calculate order parameters, correlation functions, specific heat and susceptibility. Moreover it is possible to reproduce the two observed critical temperatures.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Dirk Schmalfuß

Address for correspondence : Institut für Theoretische Physik Universität Magdeburg PF 4120, D-39016 Magdeburg, Germany

Email address : dirk.schmalfuss@physik.uni-magdeburg.de